# Information-theoretic minimax and submodular optimization algorithms for multivariate Markov chains

### v1

@@ 1 Introduction
Multivariate Markov chains on product spaces $\mathcal{X}=\mathcal{X}^{(1)}\times\ldots\times\mathcal{X}^{(d)}$ with $d\in\mathbb{N}$ arise naturally throughout stochastic modeling, Markov chain Monte Carlo (MCMC), and interacting particle systems. In high dimensions when $d$ is large, it is natural—both for analysis and for algorithm design—to approximate a complex transition matrix $P$ by a simpler model that factorizes across groups of coordinates. This paper develops an information-theoretic framework, associated structure theorems, and algorithms for selecting such factorizations and for aggregating multiple candidate Markov chains in a robust or minimax sense.
@@
This manuscript centers on the following three main threads: information projections of multivariate Markov chains, minimax information aggregation, and submodular optimization over partition. In the literature, [1] views factorization as minimizing the KL divergence between the original chain and the set of factorizable Markov chains; [2] introduces the independent projection of diffusion processes through the lens of relative entropy minimization in the space of product measures. On the topic of minimax information aggregation, [3, 4] study minimax optimization under KL divergence and $f$-divergences of probability measures, while [5] analyzes minimax excess risk as a zero-sum game between a learner and Nature. As for (robust) submodular optimization over partition, [6] and [7] propose greedy-based algorithms when the partition set function is submodular or $k$-submodular; [8] handles robust submodular optimization with bi-level optimization; [9] proposes novel algorithm with non-uniform partitions; [10] applies continuous submodular functions to address the robust budget allocation problem.
@@
Precisely, we denote $\mathcal{L}(\mathcal{X})$ to be the set of transition matrices on $\mathcal{X}$. Let $\mathcal{B}=\{P_{1},\ldots,P_{n}\}\subset\mathcal{L}(\mathcal{X})$ be a family of $\pi$-stationary transition matrices on $\mathcal{X}$ and let $\mathbf{S}=(S_{1},\ldots,S_{m})$ be a partition of $\llbracket d\rrbracket$. We consider the class of factorizable transition matrices with respect to the partition $\mathbf{S}$
@@
We interpret the minimax problem (2) in a two-person zero-sum game in Section 4, and prove that a mixed strategy Nash equilibrium always exists. This generalizes the reversiblization entropy games in [11] to the context of factorizations of multivariate Markov chains as in this paper.
@@
is orthant submodular. This result enables greedy-style algorithms with provable guarantees when designing partitions [12].
@@ 2.1 Projection and averaging of multivariate Markov chains
We consider a finite $d$-dimensional state space described by $\mathcal{X}=\mathcal{X}^{(1)}\times\ldots\times\mathcal{X}^{(d)}$. We write $\llbracket d\rrbracket:=\{1,2,\ldots,d\}$. For $S\subseteq\llbracket d\rrbracket$, we write $\mathcal{X}^{(S)}=\times_{i\in S}\mathcal{X}^{(i)}$ and $\mathcal{X}^{(-S)}=\times_{i\notin S}\mathcal{X}^{(i)}$, which are subsets of $\mathcal{X}$. We denote $\mathcal{L}(\mathcal{X})$ to be the set of transition matrices on $\mathcal{X}$, and $\mathcal{P}(\mathcal{X})=\{\pi;~\min_{x\in\mathcal{X}}\pi(x)>0,\sum_{x}\pi(x)=1\}$ to be the set of probability masses with full support on $\mathcal{X}$. We say that $P\in\mathcal{L}(\mathcal{X})$ is $\pi$-stationary with $\pi\in\mathcal{P}(\mathcal{X})$ if it satisfies $\pi=\pi P$.
We then recall the definition of the tensor product of transition matrices and probability masses, see e.g. Exercise 12.6 of [13]. Define, for $M_{l}\in\mathcal{L}(\mathcal{X}^{(l)})$, $\pi_{l}\in\mathcal{P}(\mathcal{X}^{(l)})$, $x^{l},y^{l}\in\mathcal{X}^{(l)}$ for $l\in\{i,j\},i\neq j\in\llbracket d\rrbracket,$
@@
To define the projection operations, we recall the definition of keep-$S$-in and leave-$S$-out matrices of a given transition probability matrix $P$, see Section 2.2 of [1]. For $\pi\in\mathcal{P}(\mathcal{X})$, $P\in\mathcal{L}(\mathcal{X})$, $S\subseteq\llbracket d\rrbracket$, and any $(x^{(-S)},y^{(-S)})\in\mathcal{X}^{(-S)}\times\mathcal{X}^{(-S)}$, we define the leave-$S$-out transition matrix with respect to $\pi$ to be $P_{\pi}^{(-S)}$ with entries given by
@@ 2.2 Some information-theoretic results in Markov chain theory
We first recall the Shannon entropy of a probability distribution and the entropy rate of a transition probability matrix, see Section 1 of [14]. For $\pi\in\mathcal{P}(\mathcal{X})$, its Shannon entropy is defined as
@@
We then recall the KL divergence between Markov chains (see Definition 2.1 of [1]). For given $\pi\in\mathcal{P}(\mathcal{X})$ and transition matrices $M,L\in\mathcal{L}(\mathcal{X})$, we define the KL divergence from $L$ to $M$ with respect to $\pi$ as
@@
Inspired by Theorem 2.22 of [1], we note that
@@
where the last equality comes from the fact that the averaging and projection operation commutes. ∎
@@ 2.3 Definition and examples of submodularity
We first recall the definition of a submodular function (Section 14 of [15]) and its generalization to $k$-submodularity. Given a finite nonempty ground set $U$, a set function $f:2^{U}\to\mathbb{R}$ defined on subsets of $U$ is called submodular if for all $S,T\subseteq U$,
@@
A multivariate generalization of submodularity is known as $k$-submodularity [16] where $k\in\mathbb{N}$. Let $f:(k+1)^{U}\to\mathbb{R}$ be a set function. The function $f$ is said to be $k$-submodular if
@@
We proceed to recall the definition of orthant submodularity [16]. For $\mathbf{S}=(S_{1},\ldots,S_{k}),\mathbf{T}=(T_{1},\ldots,T_{k})\in(k+1)^{U}$, let $\Delta_{e,i}f(\mathbf{S})$ be the marginal gain of adding $e$ to the $i$-th set of $\mathbf{S}$:
@@
for all $i\in\llbracket k\rrbracket$ and $\mathbf{S},\mathbf{T}\in(k+1)^{U}$ such that $\mathbf{S}\preceq\mathbf{T}$, $e\notin\mathrm{supp}(\mathbf{T})$.
@@
From Proposition 2.33 of [1], item (1) and item (2) hold. Since the map $S\mapsto H(P^{(S)})$ is submodular, the map $S\mapsto H(\overline{P}^{(S)})$ is submodular since $\overline{P}^{(S)}$ is the projection of $\overline{P}$ onto subset $S$, which proves item (3). Since
@@
we can conclude that $S\mapsto\sum_{i=1}^{n}w_{i}D_{\mathrm{KL}}^{\pi}(P_{i}\|\overline{P}^{(S)}\otimes\overline{P}^{(-S)})$ is submodular because both the map $S\mapsto H(\overline{P}^{(S)})$ and the map $S\mapsto H(\overline{P}^{(-S)})$ are submodular (by Lemma 2.1 of [12]). ∎
@@ 3 The minimax optimization problem
$\min_{A\in\mathcal{A}}\max_{P\in\mathcal{B}}\widetilde{D}_{\mathrm{KL}}^{\pi}(P\|A),$ (7)
and we can reformulate it as
@@
Suppose $A\in\mathcal{A}$ such that $\exp A(x,y)=\prod_{i=1}^{l}Q_{i}(x,y)^{c_{i}}$ for any $x,y$, we then show a Pythagorean identity based on the proof of Theorem 2.22 of [1]:
@@
We first show item (1), i.e., strong duality holds for problem (8). We shall show that the Slater’s qualification is verified (see Section 5.2.3 of [17] and Appendix A of [18]), which requires that the constraints in (8) are strictly feasible. We take any $A$ and
@@
As the strong duality in item (1) holds, by Section 5.5.2 of [17], the complementary slackness condition holds, i.e.
@@
is concave since it is the Lagrangian dual function of problem (8) (see Section 5.1.2 of [17]). ∎
@@ 4 An information-theoretic game
Inspired by the reversiblization entropy games in [11], we cast the minimax problem as a two–player zero–sum game between Nature and a probabilist. Nature chooses a transition probability matrix $P\in\mathcal{B}$, while the probabilist chooses an approximating factorizable transition matrix $Q\in\mathcal{F}=\mathcal{F}(\mathbf{S})$. The payoff from the probabilist to Nature is the KL divergence $D_{\mathrm{KL}}^{\pi}(P\|Q)$, which Nature aims to maximize while the probabilist aims to minimize.
@@
$\overline{v}=\min_{Q\in\mathcal{F}}\max_{P\in\mathcal{B}}D_{\mathrm{KL}}^{\pi}(P\|Q)=\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{\mathrm{KL}}^{\pi}(P_{i}\|\otimes_{j=1}^{m}\overline{P}^{(S_{j})}).$ (14)
The following theorem establishes the existence of a mixed-strategy Nash equilibrium (see Section 3 of [19]), which is a foundational result in game theory.
@@
    $\overline{V}(\mathbf{S},\mathcal{B})=\underline{V}(\mathbf{S},\mathcal{B})=\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{\mathrm{KL}}^{\pi}(P_{i}\|\otimes_{j=1}^{m}\overline{P}^{(S_{j})}).$
@@
We first show existence in item (1). By Proposition 3.10 of [11], we have the standard minimax inequalities $\overline{v}(\mathbf{S},\mathcal{B})\geq\overline{V}(\mathbf{S},\mathcal{B})\geq\underline{V}(\mathbf{S},\mathcal{B})$. We can also establish a lower bound for $\underline{V}$ by restricting Nature’s strategy space from all probability measures $\mathcal{P}(\mathcal{B})$ to the simplex of finite measures $\mathcal{S}_{n}$:
@@
$=\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{\mathrm{KL}}^{\pi}(P_{i}\|\otimes_{j=1}^{m}\overline{P}^{(S_{j})})=\overline{v},$
@@ 5 A projected subgradient algorithm
and then we have
@@
Inspired by Algorithm 1 of [11], we propose a projected subgradient algorithm to solve problem (15). In Algorithm 1, we conduct the projected subgradient algorithm for $t$ iterations. At each iteration, we first update the weight parameters via subgradient,
@@
which can be accomplished by existing projection algorithms onto a simplex (see e.g. [20]). Note that the subgradient algorithm is not a descent algorithm, hence the monotonicity of $h(\mathbf{w})$ among different iterations is not guaranteed, see Section 7.2 for examples.
@@
From the convexity of KL divergence $D_{\mathrm{KL}}^{\pi}(\cdot\|\cdot)$ and Equation 3.25 of [21], we have a constant $C$ such that
@@ 6 A max-min-max submodular optimization problem and a two-layer subgradient-greedy algorithm
In view of Theorem 2.6 of [12], since the map (18) is orthant submodular, then for any $\beta=\beta(\mathbf{w})\in\mathbb{R}$, if $\mathbf{S}\preceq\mathbf{V}$, we have the following monotonically non-decreasing (m1)(m-1)-submodular function:
@@
we can apply Algorithm 3 of [12] with $g$ as in (6), $c$ as in (20), and $\beta$ as in (21) to solve the problem. Furthermore, Theorem 2.11 of [12] gives the following lower bound:
@@
where $\mathbf{S}_{l}=(S_{l,1},\ldots,S_{l,m-1})$ is the final output of Algorithm 3 of [12] and $\mathbf{OPT}=\operatorname*{arg\,max}_{\mathbf{S}\preceq\mathbf{V};~|\mathrm{supp}(\mathbf{S})|\leq l}f(\mathbf{S})$.
We propose Algorithm 2 to solve problem (17). Algorithm 2 is a two-layer subgradient-greedy algorithm, which combines the outer generalized distorted greedy algorithm (Algorithm 3 of [12]) and the inner projected subgradient algorithm (Algorithm 1). Specifically, we conduct totally $l$ rounds of generalized distorted greedy algorithm: at the $i$-th round, we first fix $\mathbf{S}_{i}$ and apply the projected subgradient algorithm on fixed $\mathbf{S}_{i}$ for $K$ iterations to maximize the objective function $f(\mathbf{S}_{i},\cdot)$; we then fix $\overline{\mathbf{w}}_{i+1}=\sum_{k=1}^{K}\mathbf{w}_{i+1}^{(k)}$ and perform generalized distorted greedy algorithm to obtain $\mathbf{S}_{i+1}$. We proceed to state and prove a lower bound of Algorithm 2 in Theorem 6.3.
@@
We first refer to the proof of Theorem 2.11 of [12] and state the lower bound of the gain in the distorted greedy algorithm

### v2

@@ 1 Introduction
Multivariate Markov chains on product spaces $\mathcal{X}=\mathcal{X}^{(1)}\times\ldots\times\mathcal{X}^{(d)}$ with $d\in\mathbb{N}$ arise naturally throughout stochastic modeling, Markov chain Monte Carlo (MCMC), and interacting particle systems. In high dimensions when $d$ is large, it is natural, both for analysis and for algorithm design, to approximate a complex transition matrix $P$ by a simpler model that factorizes across groups of coordinates. This paper develops an information-theoretic framework, associated structure theorems, and algorithms for selecting such factorizations and for aggregating multiple candidate Markov chains in a robust or minimax sense.
@@
This manuscript centers on the following three main threads: information projections of multivariate Markov chains, minimax information aggregation, and submodular optimization over partition. In the literature, [choi2024geometry] views factorization as minimizing the KL divergence between the original chain and the set of factorizable Markov chains; [lacker2025independent] introduces the independent projection of diffusion processes through the lens of relative entropy minimization in the space of product measures. On the topic of minimax information aggregation, [Haussler97, Gushchin06] study minimax optimization under KL divergence and $f$-divergences of probability measures, while [HafezKolahi2022Information] analyzes minimax excess risk as a zero-sum game between a learner and Nature. As for (robust) submodular optimization over partition, [nemhauser1978analysis] and [ward2016maximizing] propose greedy-based algorithms when the partition set function is submodular or $k$-submodular; [orlin2018robust] handles robust submodular optimization with bi-level optimization; [bogunovic2017robust] proposes novel algorithm with non-uniform partitions; [staib2019robust] applies continuous submodular functions to address the robust budget allocation problem.
@@
Precisely, we denote $\mathcal{L}(\mathcal{X})$ to be the set of transition matrices on $\mathcal{X}$. Let $\mathcal{B}=\{P_{1},\ldots,P_{n}\}\subset\mathcal{L}(\mathcal{X})$ be a family of $\pi$-stationary transition matrices on $\mathcal{X}$ and let $\mathbf{S}=(S_{1},\ldots,S_{m})$ be a partition of $\llbracket d\rrbracket$, where we write $\llbracket d\rrbracket:=\{1,2,\ldots,d\}$. We consider the class of factorizable transition matrices with respect to the partition $\mathbf{S}$
@@
We interpret the minimax problem (2) in a two-person zero-sum game in Section 4, and prove that a mixed strategy Nash equilibrium always exists. This generalizes the reversiblization entropy games in [choi2023markov] to the context of factorizations of multivariate Markov chains as in this paper.
@@
is orthant submodular, where we define $\mathrm{supp}(\mathbf{S}):=\cup_{i=1}^{m-1}S_{i}.$ This result enables greedy-style algorithms with provable guarantees when designing partitions [lai2025information].
@@ 2.1 Projection and averaging of multivariate Markov chains
We consider a finite $d$-dimensional state space described by $\mathcal{X}=\mathcal{X}^{(1)}\times\ldots\times\mathcal{X}^{(d)}$. For $S\subseteq\llbracket d\rrbracket$, we write $\mathcal{X}^{(S)}=\times_{i\in S}\mathcal{X}^{(i)}$ and $\mathcal{X}^{(-S)}=\times_{i\notin S}\mathcal{X}^{(i)}$, which are subsets of $\mathcal{X}$. We denote $\mathcal{L}(\mathcal{X})$ to be the set of transition matrices on $\mathcal{X}$, and $\mathcal{P}(\mathcal{X})=\{\pi;~\min_{x\in\mathcal{X}}\pi(x)>0,\sum_{x}\pi(x)=1\}$ to be the set of probability masses with full support on $\mathcal{X}$. We say that $P\in\mathcal{L}(\mathcal{X})$ is $\pi$-stationary with $\pi\in\mathcal{P}(\mathcal{X})$ if it satisfies $\pi=\pi P$.
We then recall the definition of the tensor product of transition matrices and probability masses, see e.g. Exercise 12.6 of [levin2017markov]. Define, for $M_{l}\in\mathcal{L}(\mathcal{X}^{(l)})$, $\pi_{l}\in\mathcal{P}(\mathcal{X}^{(l)})$, $x^{l},y^{l}\in\mathcal{X}^{(l)}$ for $l\in\{i,j\},i\neq j\in\llbracket d\rrbracket,$
@@
To define the projection operations, we recall the definition of keep-$S$-in and leave-$S$-out matrices of a given transition probability matrix $P$, see Section 2.2 of [choi2024geometry]. For $\pi\in\mathcal{P}(\mathcal{X})$, $P\in\mathcal{L}(\mathcal{X})$, $S\subseteq\llbracket d\rrbracket$, and any $(x^{(-S)},y^{(-S)})\in\mathcal{X}^{(-S)}\times\mathcal{X}^{(-S)}$, we define the leave-$S$-out transition matrix with respect to $\pi$ to be $P_{\pi}^{(-S)}$ with entries given by
@@ 2.2 Some information-theoretic results in Markov chain theory
We first recall the Shannon entropy of a probability distribution and the entropy rate of a transition probability matrix, see Section 1 of [polyanskiy2025information]. For probability distribution $\pi$ on $\mathcal{X}$, its Shannon entropy is defined as
@@
We then recall the KL divergence between Markov chains (see Definition 2.1 of [choi2024geometry]). For given $\pi\in\mathcal{P}(\mathcal{X})$ and transition matrices $M,L\in\mathcal{L}(\mathcal{X})$, we define the KL divergence from $L$ to $M$ with respect to $\pi$ as
@@
Inspired by Theorem 2.22 of [choi2024geometry], we note that
@@
where the last equality comes from the fact that the averaging and projection operation commutes. As a consequence, since
$D_{\mathrm{KL}}^{\pi^{(S_{j})}}(\overline{P}^{(S_{j})}\|Q^{(S_{j})})\geq 0,$
we therefore see that
@@ 2.3 Definition and examples of submodularity
We first recall the definition of a submodular function (Section 14 of [korte2008combinatorial]) and its generalization to $k$-submodularity. Given a finite nonempty ground set $U$ as the fixed base domain, i.e. $U=\llbracket d\rrbracket$ in the context of this paper, a set function $f:2^{U}\to\mathbb{R}$ defined on subsets of $U$ is called submodular if for all $S,T\subseteq U$,
@@
A multivariate generalization of submodularity is known as $k$-submodularity [ene2022streaming] where $k\in\mathbb{N}$. Let $f:(k+1)^{U}\to\mathbb{R}$ be a set function. The function $f$ is said to be $k$-submodular if
@@
We proceed to recall the definition of orthant submodularity [ene2022streaming]. For $\mathbf{S}=(S_{1},\ldots,S_{k}),\mathbf{T}=(T_{1},\ldots,T_{k})\in(k+1)^{U}$, let $\Delta_{e,i}f(\mathbf{S})$ be the marginal gain of adding $e$ to the $i$-th set of $\mathbf{S}$:
@@
for all $i\in\llbracket k\rrbracket$ and $\mathbf{S},\mathbf{T}\in(k+1)^{U}$ such that $\mathbf{S}\preceq\mathbf{T}$, $e\notin\mathrm{supp}(\mathbf{T})$, where $\mathbf{S}\preceq\mathbf{T}$ if and only if $S_{i}\subseteq T_{i}$, $\forall i\in\llbracket k\rrbracket$.
@@
From Proposition 2.33 of [choi2024geometry], item (1) and item (2) hold. Since the map $S\mapsto H(P^{(S)})$ is submodular, the map $S\mapsto H(\overline{P}^{(S)})$ is submodular since $\overline{P}^{(S)}$ is the projection of $\overline{P}$ onto subset $S$, which proves item (3). Since
@@
we can conclude that $S\mapsto\sum_{i=1}^{n}w_{i}D_{\mathrm{KL}}^{\pi}(P_{i}\|\overline{P}^{(S)}\otimes\overline{P}^{(-S)})$ is submodular because both the map $S\mapsto H(\overline{P}^{(S)})$ and the map $S\mapsto H(\overline{P}^{(-S)})$ are submodular (by Lemma 2.1 of [lai2025information]). ∎
@@ 3 The minimax optimization problem
$\min_{A\in\mathcal{A}}\max_{P\in\mathcal{B}}\widetilde{D}_{\mathrm{KL}}^{\pi}(P\|A).$ (7)
By [candan2020chebyshev, equation (3)], we can reformulate problem (7) as
@@
Suppose $A\in\mathcal{A}$ such that $\exp A(x,y)=\prod_{i=1}^{l}Q_{i}(x,y)^{c_{i}}$ for any $x,y$, we then show a Pythagorean identity based on the proof of Theorem 2.22 of [choi2024geometry]:
@@
We first show item (1), i.e., strong duality holds for problem (8). We shall show that the Slater’s qualification is verified (see Section 5.2.3 of [boyd2004convex] and Appendix A of [beck2017first]), which requires that the constraints in (8) are strictly feasible. We take any $A$ and
@@
As the strong duality in item (1) holds, by Section 5.5.2 of [boyd2004convex], the complementary slackness condition holds, i.e.
@@
is concave since it is the Lagrangian dual function of problem (8) (see Section 5.1.2 of [boyd2004convex]). ∎
@@ 4 An information-theoretic game
Inspired by the reversiblization entropy games in [choi2023markov], we cast the minimax problem as a two–player zero–sum game between Nature and a probabilist. Nature chooses a transition probability matrix $P\in\mathcal{B}$, while the probabilist chooses an approximating factorizable transition matrix $Q\in\mathcal{F}=\mathcal{F}(\mathbf{S})$. The payoff from the probabilist to Nature is the KL divergence $D_{\mathrm{KL}}^{\pi}(P\|Q)$, which Nature aims to maximize while the probabilist aims to minimize.
@@
$\overline{v}=\min_{Q\in\mathcal{F}}\max_{P\in\mathcal{B}}D_{\mathrm{KL}}^{\pi}(P\|Q)=\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{\mathrm{KL}}^{\pi}(P_{i}\|\otimes_{j=1}^{m}\overline{P}(\mathbf{w})^{(S_{j})}).$ (14)
The following theorem establishes the existence of a mixed-strategy Nash equilibrium (see Section 3 of [osborne1994course]), which is a foundational result in game theory.
@@
    $\overline{V}(\mathbf{S},\mathcal{B})=\underline{V}(\mathbf{S},\mathcal{B})=\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{\mathrm{KL}}^{\pi}(P_{i}\|\otimes_{j=1}^{m}\overline{P}(\mathbf{w})^{(S_{j})}).$
@@
We first show existence in item (1). By Proposition 3.10 of [choi2023markov], we have the standard minimax inequalities $\overline{v}(\mathbf{S},\mathcal{B})\geq\overline{V}(\mathbf{S},\mathcal{B})\geq\underline{V}(\mathbf{S},\mathcal{B})$. We can also establish a lower bound for $\underline{V}$ by restricting Nature’s strategy space from all probability measures $\mathcal{P}(\mathcal{B})$ to the simplex of finite measures $\mathcal{S}_{n}$:
@@
$=\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{\mathrm{KL}}^{\pi}(P_{i}\|\otimes_{j=1}^{m}\overline{P}(\mathbf{w})^{(S_{j})})=\overline{v},$
@@ 5 A projected subgradient algorithm
where $|\mathcal{X}|$ denotes the cardinality of the state space $\mathcal{X}$. Then we have
@@
Inspired by Algorithm 1 of [choi2023markov], we propose a projected subgradient algorithm to solve problem (15). In Algorithm 1, we conduct the projected subgradient algorithm for $t$ iterations. At each iteration, we first update the weight parameters via subgradient,
@@
which can be accomplished by existing projection algorithms onto a simplex (see e.g. [condat2016fast]). Note that the subgradient algorithm is not a descent algorithm, hence the monotonicity of $h(\mathbf{w})$ among different iterations is not guaranteed, see Section 7.2 for examples.
@@
From the convexity of KL divergence $D_{\mathrm{KL}}^{\pi}(\cdot\|\cdot)$ and Equation 3.25 of [csiszar1972class], we have a constant $C$ such that
@@ 6 A max-min-max submodular optimization problem and a two-layer subgradient-greedy algorithm
In view of Theorem 2.6 of [lai2025information], since the map (18) is orthant submodular, then for any $\beta=\beta(\mathbf{w})\in\mathbb{R}$, if $\mathbf{S}\preceq\mathbf{V}$, we have the following monotonically non-decreasing (m1)(m-1)-submodular function:
@@
we can apply Algorithm 3 of [lai2025information] with $g$ as in (6), $c$ as in (20), and $\beta$ as in (21) to solve the problem. Furthermore, Theorem 2.11 of [lai2025information] gives the following lower bound:
@@
where $\mathbf{S}_{l}=(S_{l,1},\ldots,S_{l,m-1})$ is the final output of Algorithm 3 of [lai2025information] and $\mathbf{OPT}=\operatorname*{arg\,max}_{\mathbf{S}\preceq\mathbf{V};~|\mathrm{supp}(\mathbf{S})|\leq l}f(\mathbf{S})$.
We propose Algorithm 2 to solve problem (17). Algorithm 2 is a two-layer subgradient-greedy algorithm, which combines the outer generalized distorted greedy algorithm (Algorithm 3 of [lai2025information]) and the inner projected subgradient algorithm (Algorithm 1). Specifically, we conduct totally $l$ rounds of generalized distorted greedy algorithm: at the $i$-th round, we first fix $\mathbf{S}_{i}$ and apply the projected subgradient algorithm on fixed $\mathbf{S}_{i}$ for $K$ iterations to maximize the objective function $f(\mathbf{S}_{i},\cdot)$; we then fix $\overline{\mathbf{w}}_{i+1}=\sum_{k=1}^{K}\mathbf{w}_{i+1}^{(k)}$ and perform generalized distorted greedy algorithm to obtain $\mathbf{S}_{i+1}$. We proceed to state and prove a lower bound of Algorithm 2 in Theorem 6.3.
@@
We first refer to the proof of Theorem 2.11 of [lai2025information] and state the lower bound of the gain in the distorted greedy algorithm